OpenStudy (anonymous):

Evaluate the expression when m = 2. 11 – 3m A. 6 B. 10 C. 5 D. 12 AND Which expressions are equal to 8 when x = 6 and y = 2? Choose all answers that are correct. A. x – y + 4 B. y + x – 4 C. xy – 2 D. 5 + (x ÷ y)

OpenStudy (danjs):

replace m with 2, and calulate it

OpenStudy (danjs):

11 - 3*2

Mehek (mehek14):

\(11-3*2\\-3*2=?\)

OpenStudy (anonymous):

11-6=5?

Mehek (mehek14):

correct

Mehek (mehek14):

for the next one, put 6 for x and 2 for y then see which one is equal to 8

Mehek (mehek14):

like for A \(6-2+4=?\)

OpenStudy (anonymous):

8

Mehek (mehek14):

so that's one of the answer

OpenStudy (anonymous):

yeah the second question though needs more than one answer

Mehek (mehek14):

I know so try with the rest of the answer choices

Mehek (mehek14):

I'm just saying A is one of them

OpenStudy (anonymous):

a and d

Mehek (mehek14):

yes those are the only 2

OpenStudy (anonymous):

6/2=3+5=8

OpenStudy (anonymous):

okay thank you

Mehek (mehek14):

your welcome